# MoSE: Skill-by-Skill Mixture-of-Expert Learning for Autonomous Driving

### v1

@@ 3.2 Expert Networks and Training
Shikra-7B [4] 12.24 22.94 10.20 15.12
LLaVA1.5-7B [25] 19.30 42.06 23.16 28.17
Qwen-VL-Chat-7B [1] 18.22 26.62 22.06 22.30
MiniCPM-V-2.5-8B [38] 41.12 57.20 48.48 48.93
LLaVA1.5-13B [25] 24.54 42.41 27.90 31.61
LLaVA-NeXT-13B [27] 29.86 53.63 31.92 38.47
InternVL-V1-5-20B [8] 38.38 61.53 41.18 47.03
Qwen-VL-Max [9] 34.60 68.17 67.40 56.72
LLaVA-one [21] 38.70 51.70 49.32 46.57
@@
CODA-8B [5] 55.04 77.68 58.14 63.62
Mini-Driver-83M [41] 21.60 62.15 45.40 43.05
Mini-Drive-137M [41] 24.60 66.34 45.44 45.46
TSLVLM-7B [37] 52.84 78.33 61.62 64.26
DriveMM-8B [16] 52.94 77.76 61.84 64.18
Drive-OV-8B [16] 53.60 77.64 58.72 63.32
@@
Table 1: Performance comparison of different models. The results in gray are based on closed-source models.The row in gray are the baseline models and the row in yellow are our proposed models. The results in bold indicate the best performance within each category. Compared to models based on open-source models and data, our approach achieves SOTA performance with significantly reduced activated model size
4 Experimental Results
Dataset Settings
We conduct our main experiments on CODA dataset [5] which is designed for autonomous driving reasoning task. The dataset collects driving scenarios and extract objects which may influence ego vehicle, the final driving suggestion is generated considering all key objects. Compared to other driving reasoning dataset, CODA focus on multi-modal corner cases and consider the hierarchy data structure. It require models to generate relatively long answers containing description and analyzing of multiple objects. In this paper, we follow the official training and testing settings, and use GPT-4o for evaluation. For skill annotation, following the pipeline we introduce in Sec. 3.1.1, we obtain a skill definition as in Fig. 5.
Model Settings
We experiment the MoSE based on Qwen2-VL-2B [2] which is a small-scale VLM. We adapt a total of $14$ MoSE layers. Experimental results indicate that a relatively uniform hierarchical distribution yields better performance. Therefore, in our final configuration, we allocate $[4,5,5]$ layers to perception, prediction, and planning, respectively.
The input image size is set to $[800,450]$ to balance the image quality and GPU memory. We follow the sparse setting of general MoE that only half of the layers are equipped with MoSE. During the pretraining stage and finetuning stage, we set expert number to $4$ and use $1$ shared expert. Additionally, we also experiment with general MoE method on same backbone model (denoted as Qwen2-VL_MoE in the result table), with same number of experts and number of activated experts to our MoSE. We finetune a QwenVL2-2B under a supervised finetuning settings (denoted as Qwen2-VL_SFT). For hierarchical skill routing, we annotate $2000$ skill samples using rule-based method to pretrain the routers.
Figure 7: Visualization of expert weight distribution similarity in the final layer. The expert distribution in the MoSE model is more diverse. The similarity values are log-transformed for better clarity in reflecting differences
@@ 4 Experimental Results
Model Acc GPT Match Bleu1 Bleu2 Bleu3 Bleu4 ROUGEL CIDEr Final
MoE-StableLM 84.0 74.3 37.5 79.2 72.7 66.7 60.8 73.6 17.2 63.7
MoSE-StableLM 89.3 76.0 39.1 79.0 72.5 66.6 60.8 73.8 19.0 65.8
Table 3: QA evaluation on DriveLM-nuScences
4.1 Experiment results on CODA
Table.1 shows our results on CODA compared to other state-of-the-art models. CODA contains different subsets, general perception task lies in a comprehensive understanding of critical road key entities in driving scenarios (denoted “General”), regional perception task focus on understanding corner case objects when provided with specific bounding boxes (denoted “Regional”), and driving suggestions task aims to formulat driving advice (denoted “Suggest.”). The results are evaluated using GPT-4o, which is the same evaluation model according to CODA (denoted “Score”). The reported results contains generalist open-source and closed-source models, also specialist models which are supervised finetuned on CODA training set. As can be seen from the table, with less than 3B parameters, the proposed MoSE exceeds other SOTA methods with much less parameters. TSLVLM$+$GPT4V have the best performance but it relies on GPT for refinement.
From the results of MoE with 8 experts, the performance decline suggests that the CODA training set is insufficient to fully support the model’s training, which empirically demonstrate that general straightfoward adapting MoE encounters challenges and often results in suboptimal performance. In contrast, the proposed MoSE outperforms MoE. Further, the MoSE model with expert groups, despite having the same training data as MoE-8-1, does not suffer from performance degradation. Instead, it achieves a slight improvement, which we attribute to effective data utilization bringing by hierarchical skill routing.
Beyond the performance improvements demonstrated by the evaluation results on driving suggesting task, One thing need to be mentioned is that, the official setting of CODA uses general result and object result of context to generate final driving suggestion. In our settings, we perform single round QA without multi-turn conversation and long context.
In Fig.7 we visualize the similarity of each expert in MoE layer. During the visualization, we visualize the last MoE layer, and analyze the weights of one of the MLP layers. In a MoE framework, greater differentiation among experts enhances model performance by promoting specialization and efficient resource allocation. When experts are highly distinct, each one is able to focus on a specific subset of tasks or input distributions, leading to more refined representations and better decision-making.This differentiation helps mitigate redundancy, ensuring that experts contribute unique knowledge rather than overlapping in functionality. Additionally, a more diverse set of experts reduces competition for the same data, allowing the router to make more discriminative assignments, improving both training efficiency and generalization. As show in the Fig.7, with the hierarchical skill routing strategy, our experts have more diverse weights. For visualization of other layers, please refer to supplementary.
@@ 4.1 Experiment results on CODA
4.2 Further Analysis
@@
In Fig.8,we discussed model performance with different training data-size. During the training, we randomly sample half and quarter of general, object and driving suggestion QAs from CODA and train MoE, MoSE under the same setting with same dataset, and use the official testing set for evaluation. As the data scale increases, we observe that the performance of the MoE model peaks when trained with half data, with no further improvements as the data size continues to grow. We hypothesize that this plateau may be attributed to the way experts are allocated.
In contrast, the model trained with hierarchical skill routing exhibits a steady improvement in performance. This has been validated in both versions of the proposed MoSE. This suggests that as the dataset size increases, MoSE has the potential to achieve even better training results.
@@
Given that the proposed MoSE, compared to the original MoE, relies on a small annotated data subset, we may wonder whether this additional requirement could limit the model’s applicability to other datasets or tasks, or whether it would hinder its scalability.
@@
To address this, we design experiments to demonstrate that MoSE does not depend on large-scale additional data for pretraining but instead requires only a tiny subset. In the experiments shown in Table.2, we pretrain the model’s router using $2000$ and $3000$ skill-labeled samples, respectively. The results indicate that further scaling up the skill dataset does not lead to additional performance gains, thereby confirming that MoSE is not constrained by the need for extensive additional annotations.
4.3 Extension on other dataset
In this section, we further conduct experiment on DriveLM dataset with different backbone and skill settings. Compared to CODA, DriveLM dataset is based on nuScense dataset containing more common scenarios. Meanwhile, DriveLM considers more about driving planning and trajectory estimation, and has more diverse questions.
In the experiments, we follow a different skill settings and build a MoSE based on StableLM-1.6B[2]. For each driving scenario, we set one skill based on the scene itself, and build skill dataset using GPT-4o. In DriveLM, six-view images are used as inputs. Considering the token length limitation of the backbone model, we pair the images two by two before inputting them to reduce the number of tokens. During the experiments, we randomly select one-sixth of the original training set as the test set, while the remaining data is used for training.
Experiment results are shown in Table. 3. As can be seen in the results, the proposed MoSE has better performance with different settings in the DriveLM dataset, which once again prove the effectiveness of the proposed MoSE, and show that it can be easily extended to other tasks. (See Supplementary for more results).
5 Conclusion
@@
In this paper, we proposed a task oriented Mixture-of-Experts method named Mixture-of-Skill-Experts, which mimic human learning and reasoning process. With the proposed hierarchical skill routing strategy, the model learns from a dataset step-by-step, skill-by-skill. Empirically, MoSE demonstrates that a thoughtfully designed, skill-oriented routing mechanism can achieve superior performance in autonomous driving tasks. It effectively balances model complexity, training efficiency, and data requirements, offering a promising direction for future developments in the field. Besides, a scaling-up strategy is proposed to further boost the model’s capacity without additional data or pretraining. Further experiments show that MoSE’s hierarchical skill routing not only enhances performance with increasing data sizes but also maintains scalability by necessitating only a minimal amount of additional layers within the expert groups.
Limitation and Future work
In this work, our MoSE model is trained on a relatively small subset of available data, which means the resulting performance may still have potential to be improved. Currently, our model is designed primarily for reasoning-related tasks without direct trajectory estimation. Therefore, further experimentation is necessary to investigate whether MoSE can also benefit trajectory estimation. Additionally, integrating trajectory estimation with closed-loop evaluations, such as in simulation environments like Carla, could further enhance the practical impact of MoSE. Similar to general MoE approaches, MoSE is easily adaptable to various pre-trained models for fine-tuning on diverse downstream tasks. Such directions could be further pursued in the future.

### v2

@@ Introduction
Shikra-7B (Chen et al. 2023a) 12.24 22.94 10.20 15.12
LLaVA1.5-7B (Liu et al. 2023a) 19.30 42.06 23.16 28.17
Qwen-VL-Chat-7B (Bai et al. 2023) 18.22 26.62 22.06 22.30
MiniCPM-V-2.5-8B (Yao et al. 2024) 41.12 57.20 48.48 48.93
LLaVA1.5-13B (Liu et al. 2023a) 24.54 42.41 27.90 31.61
LLaVA-NeXT-13B (Liu et al. 2024) 29.86 53.63 31.92 38.47
InternVL-V1-5-20B (Chen et al. 2024c) 38.38 61.53 41.18 47.03
Qwen-VL-Max (Cloud 2024) 34.60 68.17 67.40 56.72
LLaVA-one (Li et al. 2024a) 38.70 51.70 49.32 46.57
@@
CODA-8B (Chen et al. 2024a) 55.04 77.68 58.14 63.62
Mini-Driver-83M (Zhang et al. 2024) 21.60 62.15 45.40 43.05
Mini-Drive-137M (Zhang et al. 2024) 24.60 66.34 45.44 45.46
TSLVLM-7B (Xue et al. ) 52.84 78.33 61.62 64.26
DriveMM-8B (Huang et al. 2024) 52.94 77.76 61.84 64.18
Drive-OV-8B (Huang et al. 2024) 53.60 77.64 58.72 63.32
@@
Table 1: Performance comparison of different models.yellow rows are our models. Bold denotes best in group. W.Exp Grp is the scaling-up MoSE model
Table 1 shows our results on CODA compared to other state-of-the-art models. CODA contains different subsets, general perception task lies in a comprehensive understanding of critical road key entities in driving scenarios (denoted “General”), regional perception tasks focus on understanding corner case objects when provided with specific bounding boxes (denoted “Regional”), and driving suggestions task aims to formulate driving advice (denoted “Suggest.”). The results are evaluated using GPT-4o, which is the same evaluation model according to CODA (denoted as “Score”). The reported results contain generalist open-source and closed-source models, as well as specialist models which are supervised finetuned on the CODA training set. As can be seen from the table, with less than 3B parameters, the proposed MoSE outperform other SOTA methods with far fewer parameters. TSLVLM++GPT4V has the best performance but it relies on GPT for refinement.
From the results of MoE with 8 experts, the performance decline suggests that the CODA training set is insufficient to fully support the model’s training, which empirically demonstrates that general straightforward application of MoE encounters challenges and often results in suboptimal performance. In contrast, the proposed MoSE outperforms MoE. Further, the MoSE model with expert groups, despite having the same training data as MoE-8-1, does not suffer from performance degradation. Instead, it achieves a slight improvement, which we attribute to effective data utilization brought by hierarchical skill routing.
Beyond the performance improvements demonstrated by the evaluation results on driving suggestion task, One thing need to be mentioned is that, the official setting of CODA uses general result and object result of context to generate the final driving suggestion. In our settings, we conduct single-round QA without multi-turn conversation and long context.
Model Score
Generalist Models
LLaVA1.5-7B (Liu et al. 2023a) 21.58
LLaVA1.6 Mistral-7B (Liu et al. 2023a) 24.09
LLaVA1.6-34B (Liu et al. 2023a) 24.94
Llama 3.2-90B (Grattafiori et al. 2024) 28.60
Qwen-2.5VL-7B (Cloud 2024) 30.63
Qwen-2.5VL-13B (Cloud 2024) 37.68
Qwen-2.5VL-72B (Cloud 2024) 37.76
Specialist Models
MoE 33.57
MoSE (Ours) 47.65
Table 2: Performance on Robo2VLM. MoSE achieves significant performance improvement compared to MoE.
@@
Experiment results on Robo2VLM
Table 1 shows our results on the Robo2VLM dataset. Robo2VLM contains multiple-choice Visual Question Answering data from human-teleoperated robot trajectorys. The score in Robo2VLM refers to the accuracy on the testing set, where each question is a 5-way multiple-choice and only one option is correct. Robo2VLM contains varying numbers of input images and covers a large diversity of scenes, which poses greater challenges for relatively small models. Compared to the suboptimal performance of the MoE baseline, our proposed skill-centric approach achieves more than 40% improvement in performance.
Figure 8: Visualization of expert weight distribution similarity in the last layer. Left:Moe, Right:MoSE. The expert distribution in the MoSE model is more diverse. The similarity values are log-transformed for better clarity in reflecting differences. Visualizations for additional layers can be found in the supplementary.
Further Analysis
@@
In Fig. 7,we discuss model performance with different training data size. During training, we randomly sample half and quarter QAs from CODA and use the official testing set for evaluation. As the data scale increases, we observe that the performance of the MoE model peaks when trained with half of the data, with no further improvements as the data size continues to grow. We hypothesize that this plateau may be attributed to the way experts are allocated. In contrast, MoSE exhibits a steady improvement in performance, with and without scaling up. This suggests that as the dataset size increases, MoSE has the potential to achieve even better training results.
@@
Given that the proposed MoSE, compared to the original MoE, relies on a small annotated data subset, we may wonder whether this additional requirement could limit the model’s applicability to other datasets or tasks, or whether it would hinder its scalability. In the experiments shown in Table 3, we pretrain the model’s router using $2000$ and $3000$ skill-labeled samples, respectively. The results indicate that further scaling up the skill dataset does not lead to additional performance gains, thereby confirming that MoSE is not constrained by the need for extensive additional annotations.
Diversity of expertise
We visualize the similarity of each expert in MoE layers. In particular, we analyze the weights of one MLP layer within the MoE modules. In a MoE framework, greater differentiation among experts enhances model performance by promoting specialization and efficient resource allocation. When experts are highly distinct, each one is able to focus on a specific subset of tasks or input distributions, leading to more refined representations and better decision-making.This helps mitigate redundancy, ensuring that experts contribute unique knowledge rather than overlapping in functionality. Additionally, a more diverse set of experts reduces competition for the same data, allowing the router to make more discriminative assignments, improving both training efficiency and generalization. As shown in the Fig. 8, with the hierarchical skill routing strategy, our experts have more diverse weights.
Conclusion
@@
In this paper, we proposed a task oriented Mixture-of-Experts method named Mixture-of-Skill-Experts, which mimics human learning and the reasoning process and learns from a dataset step-by-step, skill-by-skill. Empirically, MoSE demonstrates that a thoughtfully designed, skill-oriented routing mechanism can achieve superior performance in embodied AI tasks, such as autonomous driving and robot manipulation. Furthermore, a scaling-up strategy is proposed to further boost the model’s capacity without additional data or pretraining. Further experiments show that MoSE’s hierarchical skill routing not only enhances performance with increasing data sizes but also maintains scalability by necessitating only a small number of additional layers within the expert groups.